The cost of 2 kg of apples and 1kg of grapes on a day was found to be Rs.160. After a month, the cost of 4 kg of apples and 2 kg of grapes is Rs.300. Represent the situation algebraically

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the cost of 1 kg of apples be ‘Rs. x’
And, cost of 1 kg of grapes be ‘Rs. y’
According to the question,
the algebraic representation is 2x+y=160
And 4x+2y=300
For, 2x+y=160
or y=160−2x,
the solution table is;
x 50 60 70
y 60 40 20

For 4x+2y=300 or y = (300 – 4x)/2,
the solution table is;
x 70 80 75
y 10 -10 00

Thanks